# Toward Open Vocabulary Aerial Object Detection with CLIP-Activated Student-Teacher Learning

### v2

@@ Abstract
An increasingly massive number of remote-sensing images spurs the development of extensible object detectors that can detect objects beyond training categories without costly collecting new labeled data. In this paper, we aim to develop open-vocabulary object detection (OVD) technique in aerial images that scales up object vocabulary size beyond training data. The fundamental challenges hinder open vocabulary object detection performance: the qualities of the class-agnostic region proposals and the pseudo-labels that can generalize well to novel object categories. To simultaneously generate high-quality proposals and pseudo-labels, we propose CastDet, a CLIP-activated student-teacher open-vocabulary object Detection framework. Our end-to-end framework following the student-teacher self-learning mechanism employs the RemoteCLIP model as an extra omniscient teacher with rich knowledge. By doing so, our approach boosts not only novel object proposals but also classification. Furthermore, we devise a dynamic label queue strategy to maintain high-quality pseudo labels during batch training. We conduct extensive experiments on multiple existing aerial object detection datasets, which are set up for the OVD task. Experimental results demonstrate our CastDet achieving superior open-vocabulary detection performance, e.g., reaching 40.5% mAP, which outperforms previous methods Detic/ViLD by 23.7%/14.9% on the VisDroneZSD dataset. To our best knowledge, this is the first work to apply and develop the open-vocabulary object detection technique for aerial images.
@@ 1 Introduction
Figure 2: Challenge 2: The recall of aerial images is much lower than that of natural images. (a)(b) Aerial images from DIOR [17]. Objects in aerial images exhibit background interference. (c) Class-agnostic RPN recall statistics of novel categories in natural dataset COCO [20] and aerial dataset VisDroneZSD [23] (i.e., 48% v.s. 77%).
@@
To the best of our knowledge, this is the first work to address OVD for aerial images, and few benchmark datasets are available. We conduct extensive experiments to evaluate our method utilizing several existing aerial object detection datasets. We split the base and novel categories on these datasets, following the dataset setting in the zero-shot object detection challenge of VisDrone2023[23]. Our method achieves 40.5% mAP, surpassing Detic[47], ViLD[12] by 23.7% mAP and 14.9% mAP on VisDroneZSD dataset, showing the superiority of our proposed approach. Our contributions in this paper are summarized as follows:
@@
    Our work pioneers open vocabulary object detection in aerial imagery, aiming at addressing the fundamental challenges in earth observation image interpretation: the relatively small scale of annotated data in terms of both the number of training samples and the object categories.
@@
    We propose CastDet, a novel flexible open-vocabulary detection framework with a student-(multi)teacher interactive self-learning paradigm. We incorporate the RemoteCLIP fundamental model as an external teacher with rich knowledge to recognize novel objects in aerial imagery. Through our interactive learning and hybrid training strategy, the detection model progressively expands the object vocabulary without extra annotation efforts.
@@
    We propose a dynamic label queue for storing and incrementally updating high-quality pseudo-labels generated by the extra RemoteCLIP teacher. This strategy allows us to dynamically maintain richer and more accurate pseudo-labels from the sophisticated RemoteCLIP teacher.
@@
    We utilize several public aerial image datasets to set the open vocabulary object detection benchmark and perform extensive experiments to validate our proposal methods, achieving promising performance improvement.
@@ 3 Methodology
In this section, we firstly describe our problem setting (Sec. 3.1), followed by an overview of our CastDet framework (Sec. 3.2) following the student-teacher self-learning paradigm, and then we introduce the localization teacher as well as the reliable pseudo-bounding-boxes selection strategy (Sec. 3.3). Finally, we elaborate on the proposed dynamic pseudo-label queue to maintain the training samples with high-confidence pseudo-labels (Sec. 3.4) and the hybrid training strategy (Sec. 3.5).
@@ 3.1 Problem Description
Given a labeled detection dataset $\mathcal{L}$ with annotations on a set of base categories $\mathcal{C}_{\text{base}}$, and an unlabeled dataset $\mathcal{U}$ that may contain novel categories $\mathcal{C}_{\text{novel}}$. Our training dataset includes both labeled data and unlabled data, i.e., $\mathcal{D}_{\text{train}}=\mathcal{L}\ \cup\ \mathcal{U}=\{(I_{1},y_{1}),\cdots,(I_{n},y_{n}),I_{n+1},\cdots,I_{n+m}\}$, where $I_{i}\in\mathbb{R}^{H\times W\times 3}$ refers to the $i$-th image, and its label $y=\{(b_{i},c_{i})\}$ which consists of bounding box coordinates $b_{i}\in\mathbb{R}^{4}$ and their category $c_{i}\in\mathbb{R}^{\mathcal{C}_{\text{base}}}$. Our objective is to train a detector capable of detecting both base and novel categories, i.e., $\mathcal{C}_{\text{test}}=\mathcal{C}_{\text{base}}\cup\mathcal{C}_{\text{novel}}$, where $\mathcal{C}_{\text{base}}\cap\mathcal{C}_{\text{novel}}=\emptyset$.
@@ 3.3 Localization Teacher
Box Selection Strategies. The primary task of the localization teacher is to determine the bounding box of the objects, of which the accuracy greatly impacts the generation of reliable pseudo-labels for novel categories by the external teacher. At this stage, we place a higher emphasis on the precision of pseudo-boxes, as student-teacher learning inherently yields favorable results with fewer labels [32]. We compare different box selection strategies, as shown in Fig. 4:
@@ 3.4 Dynamic Pseudo Label Queue
The workflow of the dynamic queue comprises two steps: 1) generating high-quality pseudo-labels through the external teacher and 2) dynamically updating the pseudo-label queue and transferring data, as illustrated in Fig. 5.
Generate Pseudo Labels. Given an unlabeled image $I$ as the input of the localization teacher model, the RPN generates a set of proposals $\{p_{i}\}$. However, directly inputting these proposals to RemoteCLIP for category prediction is computationally wasteful and redundant. Therefore, we use a proposal filter to select $k$ candidates, as discussed in Sec. 3.3. Subsequently, we extract the regional features $\mathbf{f}^{P}\in\mathbb{R}^{k\times d}$ of these proposals through RoI pooling and predict their coordinates $B=\{\hat{b}_{i}\}^{k}_{i=1}$ through the class-agnostic regression branch. Finally, we obtain a set of image crops $I^{\prime}=\{x^{I}_{i}\}_{i=1}^{k}$ by cropping the corresponding areas from the image.
For an image crop $x^{I}_{i}$, we first extract its visual feature $v_{i}$ via the visual encoder of RemoteCLIP. The semantic embeddings $t_{j}$ are created as the method mentioned in Sec. 3.2. The prediction probability is performed by computing the softmax value for similarity between the visual and text semantic embeddings:
@@ 3.5 Hybrid Training
Queue Data Flow. In order to motivate the student-multi-teacher model to uncover novel objectives within the self-learning process, a specific number of images is randomly sampled from the dynamic queue, as elaborated in Sec. 3.4. Given that the localization teacher is already responsible for instructing the student in the discovery and localization of targets, we exclusively compute the classification loss for these sampled images, aiming to infuse novel knowledge into the training of the student. The objective is formatted as follows:
@@
Table 1: A summary of datasets used in our experiments. $\dagger$: DIOR serves as supplementary unlabeled training data, while NWPU-RESISC45 is employed as weakly supervised training data for the comparison experiment Detic. $\ddagger$: We crop the original images of the DOTA dataset into $\times 800$ 800 patches with an overlap of 100.
Dataset Image width # Images # Categories
total labeled unlabeled test total base novel
Det VisDroneZSD[23] 800 13067 8730 - 3337 20 16 4
NWPU VHR-10[6] $\sim$1000 800 205 - 445 10 8 2
DOTA[35] 800$\sim$4000 2806$\rightarrow$15194 10045 - 5149 15 13 2
DIOR[17] 800 23463 - 8726 - 20 - -
Cls NWPU-RESISC45[5] 256 31500 - 2100 - 45 - -
@@ 4.1 Datasets and Settings
Datasets. Due to the absence of dataset configurations specifically designed for OVD in aerial imagery currently, we follow the setup for generalized zero-shot-detection (GZSD) setting in VisDrone2023 Challenge[23] to split the base and novel categories, as shown in Table 1. We evaluate our CastDet on typical aerial datasets, including VisDroneZSD[48], DOTA[35], NWPU VHR-10[6]. We utilize a subset of DIOR[17] as supplemental unlabeled training data. For the comparison experiment Detic[47], we also need an additional classification dataset as weakly supervised information, so we provided the NWPU-RESISC45 [5] dataset specifically for Detic.
Evaluation Metrics. In the assessment of detection algorithms, the mean Average Precision (mAP), mean Average Recall (mAR) and Harmonic Mean (HM) metrics are employed as standard measures. The mAP and mAR are averaged over an Intersection Over Union (IoU) value threshold of 0.5. Following [23], we utilize the Harmonic Mean (HM) as another metric to provide a comprehensive evaluation, which is defined as the overall mAP performance of base and novel categories, i.e.,
@@
Implementation Details. We implement our method with MMDetection toolbox[3]. We employ Faster R-CNN[28] with ResNet50-C4 [15] backbone as our detection framework. The model is initialized by a pre-trained Soft Teacher [36] and R50-RemoteCLIP [21], followed by 10k iterations training with a batch size of 12 on a single A6000 GPU. Stochastic Gradient Descent (SGD) is adopted as the optimizer with a learning rate of 0.01, and the momentum and weight decay parameters are configured to 0.9 and 0.0001, respectively. To maintain high-confidence pseudo-labels, we set the RPN foreground threshold and prediction probability threshold $s_{0}$ relatively high at 0.95 and 0.8, respectively.
@@
Figure 6: Visualization of RPN foreground confidence during the training process. S, LT, ET denote student, localization teacher and external teacher, respectively.
@@ 4.2 Ablation Study
We conduct ablation studies on the VisDroneZSD dataset to thoroughly validate the effectiveness of the proposed method. These studies include different training strategies, the dynamic label queue, box selection strategies, and label fraction experiments.
Effectiveness of Hybrid Training. To demonstrate that hybrid training can effectively guide the model to discover novel categories, we compare the recall of RPN the detector under three different training mechanisms: supervised pipeline (S), closed-vocabulary student-teacher semi-supervised pipeline (S+LT) and our open-vocabulary hybrid training (S+LT+ET). For a fair comparison, our statistical recall is class-agnostic, i.e., as long as a target is proposed by RPN, it is considered to be successfully detected. As shown in Table 2, the RPN recall significantly improves by our hybrid training for novel categories compared to the semi-supervised pipeline (e.g., 69.1 vs. 47.7 mARnovel). Supervised pipeline is poor at localizing novel categories, semi-supervised pipeline performs better on only some of the novel categories, while our hybrid training approach effectively improves the detection of novel categories. Fig. 6 illustrates the visualization of RPN foreground confidence score map. While the supervised pipeline initially demonstrates the capability to discover novel category targets, its ability to identify novel categories diminishes as the model becomes adept at accurately object classification and bounding box coordinates regression. The semi-supervised pipeline can only discover categories contained in the labelled data, lacking proficiency in recognizing targets from novel categories. In contrast, Our CastDet progressively enhances its detection capability for novel categories.
Box Selection Strategy. We compare the effects of different box selection strategies. The results are shown in Table 3. The regression jittering strategy improves the performance by 4.5% mAPnovel compared to the RPN Score strategy. As depicted in Fig. 4, it exhibits a stronger correlation with both the classification score and the IoU score. So we can select more precise pseudo labels benefiting the training process.
Effectiveness of Dynamic Queue. To illustrate that the dynamic queue can dynamically obtain the high-quality categorization, we conduct experiments on whether to adopt a dynamic update strategy or whether to use a label queue. For the experimental results presented in Table 7, we can see substantial improvement achieved through applying dynamic updating or label queue techniques.
Label Fraction Experiments. To further validate that our approach is suitable for the aerial scenarios with a limited amount of labeled data, we trained our model using 34%, 50%, and 100% of the labelled data. The results are outlined in Table 5. Notably, even with a reduction to 34% of the original labeled data, the performance did not drop significantly (e.g., 39.5% mAP vs. 38.6% mAP for 100% and 34% labeled data, respectively).
Table 3: Box selection.
@@ 4.3 Comparison with the RemoteCLIP
We conducted comparison experiments with RemoteCLIP [21] and CLIP [25]. The corresponding results are presented in Table 6, revealing that CastDet exhibits an enhanced open-vocabulary classification capability compared to RemoteCLIP, i.e., from 37.6% to 47.3% mAP. Moreover, CastDet demonstrates a more significant performance improvement when directly validating RPN’s proposals, e.g., from 11.6% to 38.1% mAP. This indicates that our training approach effectively enhances novel object discovery and classification abilities. We further conduct experiments on COCO [20], and observe significant improvements with our approach on natural images as well.
@@ 4.4 Comparison with the State-of-the-Art
In Table 8, we compare the proposed method with ViLD[12] and Detic[47] on VisDroneZSD dataset. ViLD[12] distills knowledge from CLIP to train the model without requiring additional labeled data, aligning with our dataset setting. Considering the domain gap between natural and aerial images, we replace the CLIP of ViLD with RemoteCLIP to be consistent with our method for a fair comparison. Under this setting, our method surpasses ViLD by 14.9% mAP. We also compare with Detic[47], which trains the detector’s classifier on image classification data, imposing higher demands on the dataset. We provide Detic with an additional classification dataset, NWPU-RESISC45 [5]. Under this configuration, our method surpasses Detic by 23.7% mAP.
Table 8: Comparison with the state-of-the-art. $\dagger$: The best result on the zero-shot object detection leaderboard of VisDrone2023.
Method Detector #External Labeled Data Task mAP mAPbase mAPnovel HM
@@
ViLD Faster R-CNN DIOR VisDroneZSD OVD 25.6 28.5 14.2 19.0
ViLD Mask R-CNN DIOR VisDroneZSD OVD 25.4 28.7 12.1 17.0
Detic CenterNet NWPU-RESISC45 VisDroneZSD OVD 16.8 19.8 4.8 7.7
CastDet (S) Faster R-CNN DIOR VisDroneZSD OVD 35.2 33.6 41.5 37.1
CastDet (LT) Faster R-CNN DIOR VisDroneZSD OVD 40.5 39.0 46.3 42.3 $\uparrow$15.6
Figure 7: Visualization of open-vocabulary inference on VisDroneZSD dataset.

### v3

# 3OpenGVLab, Shanghai AI Laboratory

@@ Abstract
An increasingly massive number of remote-sensing images spurs the development of extensible object detectors that can detect objects beyond training categories without costly collecting new labeled data. In this paper, we aim to develop open-vocabulary object detection (OVD) technique in aerial images that scales up object vocabulary size beyond training data. The performance of OVD greatly relies on the quality of class-agnostic region proposals and pseudo-labels for novel object categories. To simultaneously generate high-quality proposals and pseudo-labels, we propose CastDet, a CLIP-activated student-teacher open-vocabulary object Detection framework. Our end-to-end framework following the student-teacher self-learning mechanism employs the RemoteCLIP model as an extra omniscient teacher with rich knowledge. By doing so, our approach boosts not only novel object proposals but also classification. Furthermore, we devise a dynamic label queue strategy to maintain high-quality pseudo labels during batch training. We conduct extensive experiments on multiple existing aerial object detection datasets, which are set up for the OVD task. Experimental results demonstrate our CastDet achieving superior open-vocabulary detection performance, e.g., reaching 46.5% mAP on VisDroneZSD novel categories, which outperforms the state-of-the-art open-vocabulary detectors by 21.0% mAP. To our best knowledge, this is the first work to apply and develop the open-vocabulary object detection technique for aerial images. The code is available at https://github.com/lizzy8587/CastDet.
@@ 1 Introduction
Figure 2: Challenge 2: The recall of aerial images is much lower than that of natural images. (a)(b) Aerial images from DIOR [18]. Objects in aerial images exhibit background interference. (c) Class-agnostic RPN recall statistics of novel categories in natural dataset COCO [21] and aerial dataset VisDroneZSD [24] (i.e., 77% v.s. 48%).
@@
To the best of our knowledge, this is the first work to address OVD for aerial images, and few benchmark datasets are available. We conduct extensive experiments to evaluate our method by comparing with the state-of-the-art open-vocabulary detectors. Additionally, we assess its performance on several existing aerial object detection datasets. We split the base and novel categories on these datasets, following the dataset setting in the zero-shot object detection challenge of VisDrone2023[24]. Our method achieves 46.5% mAPnovel, surpassing 21.0% mAPnovel on VisDroneZSD, showing the superiority of our proposed approach. Our contributions in this paper are summarized as follows:
@@
    Our work pioneers open vocabulary object detection in aerial imagery, aiming at addressing the fundamental challenges in earth observation image interpretation.
@@
    We propose CastDet, a novel flexible open-vocabulary detection framework incorporating a student-(multi)teacher self-learning paradigm. Through our hybrid training strategy, the detection model progressively expands the object vocabulary without extra annotation efforts.
@@
    We propose a dynamic label queue for storing and incrementally updating high-quality pseudo-labels, enabling us to maintain richer and more accurate labels dynamically.
@@
    We utilize several public aerial image datasets to set the open-vocabulary aerial detection benchmark and perform extensive experiments to validate our methods, achieving promising performance improvement.
@@ 3 Methodology
In this section, we firstly describe our problem setting (Sec. 3.1), followed by an overview of our CastDet framework (Sec. 3.2), and then we introduce the localization teacher as well as the reliable pseudo-bounding-boxes selection strategy (Sec. 3.3). Finally, we elaborate on the proposed dynamic pseudo label queue to maintain the training samples with high-confidence pseudo-labels (Sec. 3.4) and the hybrid training strategy (Sec. 3.5).
@@ 3.1 Problem Description
Given a labeled detection dataset $\mathcal{L}$ with annotations on a set of base categories $\mathcal{C}_{\text{base}}$, and an unlabeled dataset $\mathcal{U}$ that may contain novel categories $\mathcal{C}_{\text{novel}}$. Our training dataset includes both labeled data and unlabled data, i.e., $\mathcal{D}_{\text{train}}=\mathcal{L}\ \cup\ \mathcal{U}=\{(I_{1},y_{1}),\cdots,(I_{n},y_{n}),I_{n+1},\cdots,I_{n+m}\}$, where $I_{i}\in\mathbb{R}^{H\times W\times 3}$ refers to the $i$-th image, and its label $y_{i}=\{(b_{ik},c_{ik})\}_{k=1}^{K}$, where $K$ is the number of targets. Each target is represented by its bounding box coordinates $b_{ik}\in\mathbb{R}^{4}$ and the category label $c_{ik}\in\mathbb{R}^{\mathcal{C}_{\text{base}}}$. Our objective is to train a detector capable of detecting both base and novel categories, i.e., $\mathcal{C}_{\text{test}}=\mathcal{C}_{\text{base}}\cup\mathcal{C}_{\text{novel}}$, where $\mathcal{C}_{\text{base}}\cap\mathcal{C}_{\text{novel}}=\emptyset$.
@@ 3.3 Localization Teacher
Box Selection Strategies. The primary task of the localization teacher is to determine the bounding box of the objects, of which the accuracy greatly impacts the quality of pseudo-labels generated by the external teacher. At this stage, we place a higher emphasis on the precision of pseudo-boxes, as student-teacher learning inherently yields favorable results with fewer labels [33]. Specifically, the localization teacher first generates a set of box candidates, which are then filtered based on their confidence scores before being presented to the external teacher. We compare several indicators, as shown in Fig. 4:
@@ 3.4 Dynamic Pseudo Label Queue
The workflow of the dynamic queue comprises two steps: 1) generating high-quality pseudo-labels through the external teacher and 2) dynamically updating the pseudo label queue and transferring data, as illustrated in Fig. 5.
Generate Pseudo Labels. Given an unlabeled image $I$ as input to the localization teacher, the RPN generates a set of proposals $\{p_{i}\}$. Subsequently, we extract the regional features $\mathbf{f}^{P}\in\mathbb{R}^{k\times d}$ of these proposals through RoI pooling and predict their coordinates $B=\{\hat{b}_{i}\}^{k}_{i=1}$ through the class-agnostic regression branch. However, directly inputting these boxes to RemoteCLIP for category prediction is computationally wasteful and redundant. Therefore, we use a box filter to select $k$ candidates, as discussed in Sec. 3.3. Finally, we obtain a set of image crops $I^{\prime}=\{x^{I}_{i}\}_{i=1}^{k}$ by cropping the corresponding areas from the image.
For an image crop $x^{I}_{i}$, we first extract its visual feature $v_{i}$ via the visual encoder of RemoteCLIP. The text embeddings $t_{j}$ are generated as described in Sec. 3.2. The prediction probability is performed by computing the softmax value for similarity between the visual and text embeddings:
@@ 3.5 Hybrid Training
Queue Data Flow. In order to motivate the student-multi-teacher model to uncover novel objectives within the self-learning process, a specific number of images are randomly sampled from the dynamic label queue, as elaborated in Sec. 3.4. Given that the localization teacher is already responsible for instructing the student in the discovery and localization of targets, we exclusively compute the classification loss for these sampled images, aiming to infuse novel knowledge into the training of the student. The objective is formatted as follows:
@@ 4.1 Datasets and Settings
Datasets. Due to the absence of dataset configurations specifically designed for OVD in aerial imagery currently, we follow the setup for generalized zero-shot-detection (GZSD) setting in VisDrone2023 Challenge[24] to split the base and novel categories. We evaluate our CastDet on typical aerial datasets, including VisDroneZSD[50], DOTA[37], NWPU VHR-10[6]. We utilize a subset of DIOR[18] as supplemental unlabeled training data. For some comparison open-vocabulary detectors that require additional classification or caption data, we provide NWPU-RESISC45 [5] dataset as weak supervision. More details for datasets setting can be found in supplementary material.
Evaluation Metrics. We take the mean Average Precision (mAP), mean Average Recall (mAR) and Harmonic Mean (HM) metrics as our evaluation metrics. The mAP and mAR are averaged over an Intersection Over Union (IoU) value threshold of 0.5. Following [24], we utilize the Harmonic Mean (HM) as another metric to provide a comprehensive evaluation, which is defined as the overall mAP performance of base and novel categories, i.e.,
@@
Implementation Details. We implement our method with MMDetection toolbox[3]. We employ Faster R-CNN[29] with ResNet50-C4 [16] backbone as our detection framework. The model is initialized by a pre-trained Soft Teacher [38] and R50-RemoteCLIP [22], followed by 10k iterations training with a batch size of 12 on a single A6000 GPU. Stochastic Gradient Descent (SGD) is adopted as the optimizer with a learning rate of 0.01, and the momentum and weight decay parameters are configured to 0.9 and 0.0001, respectively. To maintain high-confidence pseudo-labels, we set the RPN foreground threshold and prediction probability threshold $p_{0}$ relatively high at 0.95 and 0.8, respectively.
@@ 4.2 Ablation Study
We conduct ablation studies on VisDroneZSD to validate the effectiveness of the proposed method. These studies include different training strategies, the dynamic label queue, box selection strategies, and label fraction experiments.
Effectiveness of Hybrid Training. To demonstrate that hybrid training can effectively guide the model to discover novel categories, we compare the recall of RPN the detector under three different training mechanisms: supervised pipeline (S), closed-vocabulary student-teacher semi-supervised pipeline (S+LT) and our open-vocabulary hybrid training (S+LT+ET). For a fair comparison, our statistical recall is class-agnostic, i.e., as long as a target is proposed by RPN, it is considered to be successfully detected. As shown in Table 1, the RPN recall significantly improves by our hybrid training for novel categories compared to the semi-supervised pipeline (e.g., 69.1 vs. 47.7 mARnovel). Supervised pipeline is poor at localizing novel categories, semi-supervised pipeline performs better on only some of the novel categories, while our hybrid training approach effectively improves the detection of novel categories.
Box Selection Strategy. We compare the effects of different box selection strategies. The results are shown in Table LABEL:tab:ab7_box_selection. The regression jittering strategy improves the performance by 4.5% mAPnovel compared to the RPN Score strategy. As depicted in Fig. 4, it exhibits a stronger correlation with both the classification score and the IoU score. Therefore, it is possible to select more precise pseudo labels, which will have a beneficial impact on the training process.
Effectiveness of Dynamic Label Queue. To illustrate that the dynamic label queue can obtain high-quality pseudo-labels, we conduct experiments on whether to adopt a dynamic update strategy or whether to use a label queue for storing pseudo-labels. In Table LABEL:tab:ab2_dynamic_queue, we can see substantial improvement achieved through applying dynamic updating or label queue techniques.
Label Fraction Experiments. To further validate that our approach is suitable for the aerial scenarios with a limited amount of labeled data, we trained our model using 34%, 50%, and 100% of the labelled data. The results are outlined in Table LABEL:tab:ab5_label_fraction. Notably, even with a reduction to 34% of the original labeled data, the performance did not drop significantly (e.g., 39.5% mAP vs. 38.6% mAP for 100% and 34% labeled data, respectively).
Table 2: Ablation experiments results. (a) Comparison on box selection strategies. (b) Evaluation on other remote sensing datasets. (c) Ablation study on fraction of labeled data. (d) Evaluation on CLIPs (e.g., RemoteCLIP, CLIP) with ground-truth (GT) or RPN pseudo boxes. (e) Ablation study on dynamic label queue.
@@ 4.3 Comparison with the RemoteCLIP
We conducted comparison experiments with RemoteCLIP [22] and CLIP [26]. The corresponding results are presented in Table LABEL:tab:com2clip, revealing that CastDet exhibits an enhanced open-vocabulary classification capability compared to RemoteCLIP, i.e., from 37.6% to 47.3% mAP. Moreover, CastDet demonstrates a more significant performance improvement when validating RPN’s proposals, e.g., from 11.6% to 38.1% mAP. This indicates that our training approach effectively enhances novel object discovery and classification abilities. We further conduct experiments on COCO [21], and observe significant improvements with our approach on natural images as well.
@@ 4.4 Comparison with the State-of-the-Art
In OVD, auxiliary weak supervision may be used, such as: 1) using both base and novel categories ($\mathcal{T}_{\mathrm{novel}}$) with unlabeled images for knowledge distillation or pseudo-labeling, such as ViLD [13] and OV-DETR [43], aligning with our setting; and 2) using classification or caption data ($\mathcal{T}_{\mathrm{cls}}$) that may contain novel concepts, such as Detic [49], etc [7, 23, 19, 35], therefore we provide an additional classification dataset NWPU-RESISC45 [5] for these methods. We compare the proposed method with above SOTA open-vocabulary detectors on VisDroneZSD dataset, as shown in Table 3. Our method outperforms the previous SOTA by 21.0% mAPnovel. To further evaluate the open-vocabulary generalization capability of our method, we construct a large concept pool for self-training without requiring novel classes to be pre-known. The results in Table 3 show comparable performance. The visualization results can be found in supplementary material.
Table 3: Comparison with the state-of-the-art detectors. $\mathcal{T}_{\mathrm{novel}}$ indicates whether the novel classes need to be pre-known. $\mathcal{T}_{\mathrm{cls}}$ denotes whether extra classification or caption datasets are required during training. $\dagger$: The best result on the zero-shot object detection leaderboard of VisDrone2023. $\ddagger$: The results of our own implementation, we replace the CLIP of ViLD with RemoteCLIP.
Method Backbone External Labeled Data Task $\mathcal{T}_{\mathrm{novel}}$ $\mathcal{T}_{\mathrm{cls}}$ mAP mAPbase mAPnovel HM
@@
ViLD [13] RN50 DIOR VisDroneZSD OVD 25.6 28.5 14.2 19.0
OV-DETR [43] RN50 DIOR VisDroneZSD OVD 25.6 25.6 25.5 25.6
Detic [49] RN50 NWPU-RESISC45 VisDroneZSD OVD 16.8 19.8 4.8 7.7
GroundingDINO [23] RN50 NWPU-RESISC45 VisDroneZSD OVD 33.4 40.5 0.3 6.1
GLIP [19] Swin-T NWPU-RESISC45 VisDroneZSD OVD 33.6 41.0 5.4 9.5
BARON [35] RN50 NWPU-RESISC45 VisDroneZSD OVD 27.4 29.4 19.5 23.5
YOLO-World [7] YOLOv8-M NWPU-RESISC45 VisDroneZSD OVD 32.9 39.1 8.5 13.9
CastDet (Ours) RN50 DIOR VisDroneZSD OVD 38.5 36.5 46.5 40.9
CastDet (Ours) RN50 DIOR VisDroneZSD OVD 40.5 39.0 46.3 42.3